# OCR Hinders RAG: Evaluating the Cascading Impact of OCR on Retrieval-Augmented Generation

### v1

@@ 3.3 Data perturbation with OCR noise
Despite advancements in OCR, real-world applications often encounter document types beyond the training corpus of OCR models, leading to low-quality data extraction. Additionally, the inherent non-uniform representation of document elements further introduces noise, affecting the performance of each component in RAG systems. In this paper, we primarily focus on two primary types of OCR noises: Semantic Noise and Formatting Noise. To systematically explore these noises, we start from errors in current OCR results and generate perturbed data with different noise levels based on the ground truth structured data. We will now delve into the details of each type.
Semantic Noise results from OCR prediction errors, including typos, misrecognized formula symbols, as well as content and structural errors in tables. This type of noise impacts the semantics of parsed content, deviating retrievers and LLMs from integrating correct information related to user queries. As its appearance is irregular, we introduce this type of noise by providing examples and adopting GPT-4o to mimic it. Specifically, we begin with gathering real examples from current OCR results. We divide them into three categories according to their edit distance to the ground truth, and prompt GPT-4o to modify ground truth structured data in three levels based on these examples. The perturbations are across all document elements, including plain text, table, and formula. As illustrated in Fig. 3, minor perturbations on the table include typos, while moderate perturbations disrupt structure by adding or misaligning rows and columns, which makes it more complicated for retrieval and LLM’s perception. Severe perturbations produce tables with control character omission, which cannot be rendered correctly, simulating OCR recognition breakdowns. For plain text and formula, misspelling, visually similar characters, and truncation are introduced in different proportions to produce three different levels of noise.
Figure 3: Illustration of different levels of Semantic Noise on plain text, equation, and table, which are all perturbed based on existing OCR results.
[…]
@@
Qwen2-VL-72B [33] 0.2564 53.16 72.97 26.72 41.23 23.45 35.91
InternVL2-Llama3-76B [5] 0.4450 42.43 57.51 20.74 32.89 20.58 31.23
Table 2: Evaluation of various OCR solutions and their impacts on RAG systems. We report the retrieval performance using top-1 (LCS@1) and top-5 (LCS@5) retrieved chunks. Overall performance is presented with top-1 (EM@1, F1@1) and top-5 (EM@5, F1@5) retrieved chunks. Bold indicates the best performance, and underline indicates the second-best performance.
@@ 4.1 Experimental settings
We evaluate the impact of OCR on RAG systems in three ways: retrieval performance, generation performance, and overall system performance. For the retrieval stage, we utilize knowledge bases derived from the same domains of user queries. During the generation stage, we provide the page where the question is derived from for LLMs to generate the response. In the overall evaluation, retrievers retrieve the relevant chunks from the knowledge base in the same domain as the question, and LLMs generate responses based on these chunks. In the overall evaluation, we provide the top-1 matched chunk for generation unless otherwise stated. The default chunk size is $1024$ with no overlap.
Metrics. To evaluate the quality of OCR results, we calculate the edit distance between each page of OCR results and the ground truth structured data and report the average values. For assessing retrieval performance, as results of different OCRs often include various extraneous characters, discriminating whether the evidence exactly appears in the retrieved contents is not fair. Following [14], we employ Longest Common Subsequence (LCS) [26] to measure evidence inclusion in retrieved content. For the generation stage, we employ the F1-score and Exact-Match metrics to measure the accuracy of LLMs’ responses.
Retrievers. We consider two primary retrievers: (1) BGE-M3 [4], a recent SOTA dense retriever within its size category. (2) BM-25 [28, 31] a lightweight sparse retriever ranking document based on the query term frequency.
LLMs. We employ two representative open-source LLMs: Qwen2 (Qwen2-7B-Instruct) [33] and Llama-3.1 (Llama3.1-8B-Instruct) [8] and the proprietary LLM: GPT-4o (gpt-4o-2024-08-06). A standard prompt template is used to format responses consistently across all LLMs (see Appendix Sec. I). All open-source models are downloaded from Huggingface  [^6] https://huggingface.co/, with inference conducted on 8 NVIDIA A100 GPUs.
@@ 4.2 Benchmarking current OCR solutions
In this section, we evaluate the suitability of current OCR solutions for real-world RAG applications by conducting comprehensive experiments with our OHRBench. We involve several representative OCR solutions including (1) Pipeline-based OCR, such as MinerU [32] and Marker [25], (2) End-to-end OCR, including GOT [34] and Nougat [2], and (3) Vision-Language Models, specifically Qwen2-VL-72B [33] and InternVL2-Llama3-76B [5]. For GOT, we employ its format OCR mode to output structured data. For Qwen2-VL-72B and InternVL2-Llama3-76B, we prompt them to produce formulas and tables in LaTeX format, with the prompt template available in Appendix Sec. I. We use PyMuPDF to convert the original PDF page to an image and use the default input resolution of each OCR solution. To ensure fair edit distance calculations, tables from different OCR results are converted into LaTeX format. The retrievers used are BGE-M3 and BM25, while the LLMs are LLama-3.1-8B-Instruct and Qwen2-7B-Instruct. All metrics are averaged across domains and combinations of retrievers and LLMs. Details of experimental results are available at Appendix Sec. III.1.
Through the comparison presented in Tab. 2, we derive several key conclusions about the performance of these OCR solutions and their corresponding impacts on RAG systems, as follows: (1) Pipeline-based OCR demonstrates the best performance. Employing Marker achieves the best retrieval performance across all OCR solutions, while MinerU dominates the generation and overall evaluation. This is due to a large number of complex layouts in our document datasets, and pipeline-based OCR can perform fine-grained local recognition based on the layout detection results, thereby a large number of complex layouts in our document datasets, and pipeline-based OCR can perform fine-grained local recognition based on the layout detection results, better-supporting issues such as reading orders and dense text. (2) The best solutions of end-to-end OCR and VLMs can achieve comparable performance. GOT and Qwen2-VL-72B come out on top in end-to-end OCR and VLMs, respectively, with Qwen2-VL-72B being the closest to the performance of pipeline-based OCR and achieving sub-optimal results in LCS@1. The main reason why they do not perform as well as pipeline-based OCR is that they have difficulty handling complex layouts, especially for Newspaper, where all end-to-end OCR and VLMs have a serious performance degradation. (See Appendix Sec. III.1 for detailed results on each domain) Meanwhile, they also struggle with out-of-training data distribution and hallucination issues. Nougat is designed for academic papers and achieves good performance, but it is difficult to generalize to other fields. InternVL2-Llama3-76B, on the other hand, struggles with hallucinations and generates a large number of repetitive and nonsensical characters when handling images with dense text. (3) All OCR solutions suffer performance loss. Even the best solutions show a decrease of 1.9 in EM@1 and 2.93 F1@1 in the overall evaluation, with greater losses in the retrieval and generation stages.
@@
In summary, current OCR solutions struggle to ensure robustness and effectiveness across diverse real-world RAG application scenarios. Moreover, the OCR metric of edit distance does not always correlate with RAG performance. For example, although MinerU and Qwen2-VL-72B exhibit lower edit distances compared to Marker, they do not consistently achieve better performance across all metrics. This discrepancy may be attributed to the varying types of OCR noise introduced by different solutions. To further investigate, we systematically explore the impact of these OCR noise types on RAG in Sec. 4.3.
@@ 4.3 In-depth analysis of OCR noise’s impact on RAG
Figure 4: Impact of Semantic Noise ([S] dashed lines) and Formatting Noise ([F] solid lines) on RAG components. The horizontal axis denotes the edit distance, where higher values indicate increased OCR noise. We report LCS@1 and F1-score for each evidence source, including text (first column), table (second column), formula (third column), and all sources combined (last column).
In this section, we conduct an in-depth analysis of the impact of Semantic Noise and Formatting Noise on RAG systems, with our perturbed structured data derived from ground truth. We employ edit distance to quantify these perturbations, with a value of 00 representing ground truth structured data.
@@ 4.3.1 Fine-grained impact on retrieval and generation
Semantic Noise significantly influences both retrieval and generation phases. As illustrated in Fig. 4, increasing perturbation levels from mild to severe results in nearly a $50\%$ performance decline for most retrievers and LLMs. In the retrieval stage, the sparse retriever BM25 initially outperforms the dense retriever BGE-M3 on ground truth structured data. However, as perturbation intensifies, BM25’s performance eventually falls below BGE-M3, indicating its vulnerability to Semantic Noise. Notably, for table-related questions, BGE-M3 shows a 10% improvement with mild perturbation compared to the ground truth structured data. This may relate to mild perturbation involving only plain text and formulas, leaving tables unaltered, thus minimizing the impact on table-related queries. In the generation phase, both proprietary and open-source LLMs struggle with Semantic Noise, among which performance on table questions decreases the most. For instance, GPT-4o experiences a 50% drop in F1-score on table-related questions, indicating the importance of correctly parsing table structure.
Formatting Noise primarily affects questions related to multimodal elements. While performance on plain text queries shows slight variation, multimodal queries, especially those involving formulas, experience significant degradation. The maximum performance drops are 16.8% for BM25 and 19.4% for Qwen2-7B in retrieval and generation, respectively. Table questions see declines of 9.9% (BM25) and 18.4% (Llama3.1-8B), respectively. GPT-4o is relatively resilient, with only 6% reduction in performance for both formula and table questions, indicating that more advanced LLMs can effectively handle Formatting Noise.
@@ 4.3.2 Impact on end-to-end evaluation
Semantic Noise consistently demonstrate a strong impact, while Formatting Noise affects specific retrievers and LLMs differently. All retrievers and LLMs struggle with Semantic Noise, leading to consistent performance drops, particularly in multimodal questions involving formulas and tables. In contrast, the effect of Formatting Noise is more variable, depending on the specific bottlenecks within the system. Despite a decrease in retrieval performance with increasing perturbations, when employing Llama3.1-8B for generation, the overall performance shows a slight change due to its limited information integration capabilities. Conversely, overall performance using GPT-4o is more sensitive to changes in the retrieval stage. For table-related questions, improvements in BGE-M3’s performance under mild perturbations lead to enhanced results with GPT-4o. However, in formula-related questions, performance deteriorates to the level of Qwen2-7B as the retrievers’ performance decreases.
@@
Figure 5: Performance of retrieval, generation and end-to-end with different table format. We only report the results of table-related questions. Other experimental results can be found in Appendix Sec. III.1.
@@ 4.3.3 Impact of table format
In addition to perturbations, we investigate the influence of different table formats, as a kind of Formatting Noise. As shown in Fig. 5, HTML tables show inferior performance during retrieval compared to Markdown and LaTeX formats, less than half the performance of Markdown tables. Markdown and LaTeX formats perform similarly, with BGE-M3 demonstrating a better understanding of Markdown. During the generation phase, the impact of table formats varies among LLMs. For example, Qwen2-7B performs better with HTML tables, whereas GPT-4o maintains consistent performance across all formats, with a maximum F1-score variation of only $2.1$. In end-to-end evaluations, affected by low retrieval performance, HTML tables result in the lowest across nearly all retriever and LLM combinations, while the combination of GPT-4o and BM25 using Markdown achieves the best performance.
4.4 Potential of employing Vision-Language Models in RAG systems.
Model Context Input Generation
EM F1
GPT-4o GT Text 50.60 70.66
OCR Text 45.62 64.67
Image 39.26 56.28
Image+OCR Text 49.97 70.05
Qwen2-VL-7B GT Text 48.70 66.58
OCR Text 43.19 61.04
Image 44.09 59.63
Image+OCR Text 48.13 63.48
InternVL2-8B GT Text 44.72 56.12
OCR Text 41.22 52.53
Image 31.10 51.16
Image+OCR Text 38.31 56.43
Table 3: Performance of employing VLMs for generation. For context input, GT Text represents ground truth structured data, OCR Text represents the OCR processed data from MinerU, and Image refers to the PDF image containing the clue.
Vision-Language Models (VLMs) have emerged as a promising solution for document understanding tasks without relying on OCR [23, 11, 6]. In this section, we evaluate the efficacy of employing VLMs in the generation stage of RAG systems. We evaluate four types of context inputs: ground truth text, OCR text, image, and a combination of image and OCR text, using MinerU for OCR due to its superior performance in Tab. 2. We keep the prompt templates consistent, with slight input format adjustments (see Appendix Sec. I). Results are shown in Tab. 3. (1) The performance of image-only input is still lagging behind inputting OCR results. For GPT-4o and InternVL2-8B, using images alone yields lower EM-scores and F1-scores compared to OCR text. Although Qwen2-VL-7B achieves comparable performance with image input, it is still inferior to the performance using ground truth text. This indicates that current VLMs are not yet capable of fully replacing OCR-based RAG systems. (2) Combining multimodal inputs enhances VLM performance. When both image and OCR text are fed, GPT-4o, Qwen2-VL-7B, and InternVL2-8B show consistent performance improvements, with F1-score increases of 24.5%, 6.5% and 10.3% respectively, compared to image-only input. Notably, the EM value of GPT-4o and Qwen2-VL-7B, and the F1-score of GPT-4o and Internvl2-8B, achieve performance comparable to ground truth text. This suggests that the detailed text information compensates for VLMs’ challenges with complex documents that may be beyond their training corpus. These findings highlight that employing OCR results is orthogonal to employing VLMs, facilitating future research to explore VLMs’ potential in RAG systems.

### v2

@@ 3.3 Data perturbation with OCR noise
Despite advancements in OCR, real-world applications often encounter document types beyond the training corpus of OCR models, leading to low-quality data extraction. Additionally, the different structured representations of document elements further introduces noise, impacting RAG performance. In this paper, we focus on two key types of OCR noises: Semantic Noise and Formatting Noise. To quantitatively analyze their effects, we start from errors in current OCR results and generate perturbed data with different noise levels. We will now delve into the details of each type.
Semantic Noise results from OCR prediction errors that impact the semantics of parsed content, deviating retrievers and LLMs from integrating correct information related to user queries. To systematically capture realistic Semantic Noise, we include diverse perturbation to document images and utilize multiple OCR solutions to perform OCR on these document images, capturing a wide range of real-world Semantic Noise as much as possible. We begin with collecting naturally distorted documents and identifying common degradation patterns, such as background artifacts, watermarks, and structural distortions (e.g., dilation and erosion). Then, we extend from [3], where its method has been shown to be effective for simulating naturally distorted documents, we refine perturbation strategies through an iterative, cross-validated process involving multiple annotators. One annotator adjusts distortion parameters and applies them to document images, while a another annotator, who is unaware of the applied modifications, distinguish which document appears artificially altered. This refinement continues until the perturbations become indistinguishable from real samples. Through this process, we identify 8 effective perturbation types that balance realism and intensity. Details and examples are provided in Appendix Sec. III.2. By varying the number and type of perturbations, we generate 3 distinct datasets with controlled Semantic Noise levels. We then choose MinerU, GOT, and Qwen2.5-VL to curate 9 perturbed data with diverse appearance of Semantic Noise, enabling a systematic evaluation of its impact on RAG performance.
Formatting Noise stems from stylistic commands, such as white space characters for beautifying formulas and bold and italic commands for better readability, and inconsistencies in structured data representations across Markdown, LaTeX, and HTML. Although irrelevant to semantics, this noise complicates information integration for both retrievers and LLMs. To assess the impact of Formatting Noise on RAG, we identify common OCR-induced formatting inconsistencies and develop heuristic rules to introduce controlled perturbations through additions, removals, and format conversions. A detailed list of perturbation rules is in Appendix Sec. III.1. By applying these modifications at varying proportions in ground truth structured data, we create three datasets with different degrees of Formatting Noise. Additionally, we evaluate RAG performance under different structured data formats, comparing retrieval and reasoning consistency across Markdown, LaTeX, and HTML representations.
@@
Qwen2.5-VL-72B [2] 0.18 74.6 59.8 59.7 38.2 5.3 59.2 44.4 42.1 31.8 27.0 11.6 37.5 40.6 31.1 26.1 19.0 8.8 31.1
InternVL2.5-78B [6] 0.28 68.2 57.7 55.3 45.1 2.7 55.8 41.8 41.8 29.0 33.6 3.3 35.8 38.2 31.0 23.3 22.9 3.1 29.6
Table 2: Evaluation of various OCR solutions and their impacts on RAG systems. The OCR performance is reported using edit distance (E.D.). We report the generalized LCS or F1 of five types of evidence sources, including plain text (TXT), table (TAB), formula (FOR), chart (CHA), and reading order (RO). Bold indicates the best performance, and underline indicates the second-best performance.
@@ 4.1 Experimental settings
We evaluate the impact of OCR on RAG systems in three ways: retrieval performance, generation performance, and overall system performance. For the retrieval stage, we utilize knowledge bases derived from the same domains of user queries and retrieve the top-2 matched chunk. During the generation stage, we provide the page where the question is derived from for LLMs to generate the response. In the overall evaluation, retrievers retrieve the relevant chunks from the knowledge base in the same domain as the question, and LLMs generate responses based on these chunks. In the overall evaluation, we provide the top-2 matched chunk for generation unless otherwise stated. The default chunk size is $1024$ with no overlap.
Metrics. To evaluate the quality of OCR results, we calculate the edit distance between each page of OCR results and the ground truth structured data and report the average values. For assessing retrieval performance, as results of different OCRs often include various extraneous characters, discriminating whether the evidence exactly appears in the retrieved contents is not fair. Following [18], we employ Longest Common Subsequence (LCS) [33] to measure evidence inclusion in retrieved content. For the generation stage, we employ the F1-score metric to measure the accuracy of LLMs’ responses.
Retrievers. We consider two primary retrievers: (1) BGE-M3 [5], a recent SOTA dense retriever within its size category. (2) BM-25 [35, 39] is a lightweight sparse retriever ranking document based on the query term frequency.
LLMs. We employ three representative open-source LLMs: Qwen2 (Qwen2-7B-Instruct and Qwen2-72B-Instruct) [42] and Llama-3.1 (Llama3.1-8B-Instruct) [11]. A standard prompt template is used to format responses consistently across all LLMs (see Appendix Sec. I). All open-source models are downloaded from Huggingface  [^3] https://huggingface.co/, with inference conducted on 8 NVIDIA A100 GPUs.
@@ 4.2 Benchmarking current OCR solutions
In this section, we evaluate the suitability of current OCR solutions for real-world RAG applications by conducting comprehensive experiments with our OHRBench. We involve several representative OCR solutions including (1) Pipeline-based OCR, such as MinerU [41] and Marker [32], (2) End-to-end OCR, including GOT [43] and Nougat [3], and (3) Vision-Language Models, specifically Qwen2.5-VL-72B [38] and InternVL2.5-78B [6]. For GOT, we employ its format OCR mode to output structured data. For Qwen2.5-VL-72B and InternVL2.5-78B, we prompt them to produce formulas, tables, and charts in LaTeX format, with the prompt template available in Appendix Sec. I. The retrievers used are BGE-M3 and BM25, while the LLMs are LLama-3.1-8B-Instruct and Qwen2-7B-Instruct. All metrics are averaged across domains and combinations of retrievers and LLMs. Details of experimental results are available in Appendix Sec. IV.1.
Through the comparison presented in Tab. 2, we derive several key conclusions about the performance of these OCR solutions and their corresponding impacts on RAG systems, as follows: (1) VLMs for OCR achieve the best overall performance. Among all OCR solutions, Qwen2.5-VL-72B consistently outperforms others across all three evaluation stages. Its superiority stems from its ability to handle structured data more effectively than both pipeline-based and end-to-end OCR methods. Despite claims that GOT can parse charts, its performance remains subpar. Similarly, MinerU, Marker, and Nougat fail to produce comparable results due to their inability in parsing chart. For plain text questions, although its poor performance on high-resolution documents (see newspaper in APPendix Sec. IV.1, Qwen2.5-VL-72B performs comparably to pipeline-based OCR. Our manual review suggests that its strong language decoder enhances robustness against historical and distorted documents, a capability lacking in pipeline-based OCR. (2) Reading order is challenging for VLMs and End-to-end OCR. Despite their strong semantic understanding capabilities offered by language decoders, both VLM and end-to-end OCR struggle with merging paragraphs correctly, reflected in F1-score of just 8.8 in overall evaluation. In contrast, pipeline-based OCR, though lacking semantic understanding, achieves performance close to ground truth (GT) using rule-based strategies. However, GT itself performs poorly, likely because reading-order questions require integrating information across multiple paragraphs, posing challenges for current RAG systems [37]. (3) All OCR solutions exhibit performance degradation. Even the best solutions experience a 14% (5 F1-score) drop in the overall stage evaluation, with greater losses in the retrieval and generation stages. This indicates that our OHRBench presents significant challenges for both OCR solutions and RAG systems.
@@
In summary, current OCR solutions struggle to maintain robustness and effectiveness across diverse real-world RAG application scenarios. Additionally, standard OCR metrics like edit distance do not always align with RAG performance. For example, while MinerU and Qwen2.5-VL-72B exhibit lower edit distances compared to Marker, they do not consistently achieve better performance across all metrics. This discrepancy may be attributed to the varying types of OCR noise introduced by different solutions. To further investigate, we systematically explore the impact of these OCR noise types on RAG in Sec. 4.3.
@@ 4.3 In-depth analysis of OCR noise’s impact on RAG
Figure 3: Impact of Semantic Noise ([S] dashed lines) and Formatting Noise ([F] solid lines) on RAG components. The horizontal axis denotes the ratio $r_{\text{noise}}$, where higher values indicate greater OCR-induced noise. We report LCS and F1-score for each evidence source: text (first column), the average score for multimodal elements (tables, formulas, and charts, second column), reading order (third column), and all sources combined (last column).
In this section, we conduct an in-depth analysis of the impact of Semantic Noise and Formatting Noise on RAG systems, using perturbed structured data with varying levels of perturbations. For each type of OCR noise, we introduce three noise levels—mild, moderate, and severe—to systematically assess their effects. Since our findings in Sec. 4.2 indicate that edit distance fails to accurately capture the degree of OCR noise, we instead define the ratio $r_{\text{noise}}$, representing the proportion of Q&A pairs affected by OCR noise as a measure. Specifically, for each Q&A pair, we compute the LCS between its evidence context and the corresponding perturbed structured data. If LCS exceeds 0.95, the Q&A pair is considered as unaffected, otherwise, it is affected. We then use the ratio ($r_{\text{noise}}$) to quantify the degree of perturbations, with 0 representing to ground truth structured data and values approaching 1 indicating greater perturbation. This approach allows us to align the perturbation levels of Formatting Noise and Semantic noise for fair comparisons. Additionally, for Formatting Noise, we evaluate its retrieval performance with modified LCS calculations by excluding stylistic commanding introduced during perturbation, ensuring a fair assessment of retrieval accuracy. For each degree of Semantic Noise, we report the average RAG performance using three different OCR results, including MinerU, GOT, and Qwen2.5-VL-72B.
@@ 4.3.1 Fine-grained impact on retrieval and generation
Semantic Noise significantly influences both retrieval and generation phases. As illustrated in Fig. 3, increasing Semantic Noise from mild ($r_{\text{noise}}=0$) to severe ($r_{\text{noise}}>0.6$) results in nearly a $50\%$ performance decline for most retrievers and LLMs. In the retrieval stage, both the sparse retriever BM25 and the dense retriever BGE-M3 suffer consistent performance declines across all types of questions, suggesting that dense retrieval’s stronger comprehension does not provide robustness against Semantic Noise. In the generation phase, all LLMs struggle with Semantic Noise, among which performance on reading-order decreases the most. Interestingly, although the way we introduce Semantic Noise should primarily affect text recognition, questions related to multimodal elements (tables, formulas and charts) degrade even further, highlighting the challenges in parsing, understanding and reasoning over multimodal document data.
Formatting Noise primarily affects multimodal questions. While performance on plain text queries and reading-order-related questions remain largely unaffected, retrieval and generation performance drops more severe for multimodal queries. The maximum performance losses reach 12.7% for BGE-M3 and 9.1% for Llama3.1-8B in retrieval and generation, respectively. In addition, larger LLMs exhibit greater robustness, with only a 7% performance reduction on multimodal questions, indicating that more advanced LLMs can effectively handle Formatting Noise.
@@ 4.3.2 Impact on end-to-end evaluation
Semantic Noise consistently demonstrate a strong impact, while Formatting Noise affects specific retrievers and LLMs differently. Semantic Noise consistently degrades performance across all combination of retrievers and LLMs, particularly on multimodal questions involving tables, formulas, and charts. In contrast, the effect of Formatting Noise is more variable when using smaller LLMs, Llama3.1-8B and Qwen2-7B, for generation, despite greatly reduced retrieval accuracy, the overall performance shows a slight change due to their limited information integration capabilities. Conversely, using a larger LLM, Qwen2-72B, is highly sensitive to retrieval performance. While its generation performance shows light changes, the overall performance decreases more, especially on questions related to multimodal elements.
@@
Figure 4: Performance of retrieval, generation and end-to-end with different table format. We only report the results of table-related questions.
@@ 4.3.3 Impact of table format
In addition to perturbations, we investigate the influence of different table formats as a kind of Formatting Noise. As shown in Fig. 4, HTML tables show inferior performance during retrieval compared to the Markdown and LaTeX formats. Markdown and LaTeX formats perform similarly, with BGE-M3 demonstrating a better understanding of Markdown. In the generation phase, HTML and LaTeX showed similar performance across all models, but the Markdown format performed worse due to the lack of support for merging cells. In end-to-end evaluations affected by low retrieval performance, using HTML tables is comparatively worse, while the combination of Qwen2-72B and BGE-M3 using Markdown achieves the best performance.
4.3.4 Error analysis
We further conduct error analysis to understand the bottlenecks of the RAG system with OCR noise in a quality approach. Specifically, we calculate the distribution of OCR errors and RAG errors when using ground truth structured data and perturbed data with severe Formatting Noise and Semantic Noise. Our evaluation employs BGE-M3 as the retriever and assesses error distributions using two generators: Qwen2-7B and Qwen2-72B. We use the same strategies to identify the proportion of OCR errors as in Sec. 4.3. The distribution of these errors is illustrated in Fig. 5. It indicates that when the proportion of OCR errors is similar for Semantic Noise and Formatting Noise (66% vs. 61%), it performs better with Formatting Noise. Of these, about half of the correct responses when using Qwen2-7B as a generator were done by the model with incorrect OCR results. Meanwhile, larger LLMs are more robust to OCR noise. Compared to Qwen2-7B, Qwen-72B has a 9.4% and 3.5% higher percentage of samples with OCR errors but ultimately correct in Formatting Noise and Semantic Noise, respectively.
Figure 5: Analysis of answer correctness distribution in Q&A pairs, using different perturbed data and LLMs. It reveals that larger LLMs are more robust to OCR noise. [S] and [F] denote perturbed data with severe Semantic Noise and Formatting Noise, respectively. [GT] represents ground truth perturbed data.